# Rethinking Pan-sharpening: Principled Design, Unified Training, and a Universal Loss Surpass Brute-Force Scaling

### v1

@@ 5 Conclusion
In this paper, we challenged the prevailing “bigger is better” paradigm in pan-sharpening by focusing on efficiency, generalization, and principled engineering. We introduced PanTiny, a lightweight and highly efficient model, and demonstrated that its carefully considered architecture achieves a superior balance of performance and computational cost compared to larger, more complex SOTA models. Our most significant contribution is the novel “all-in-one” training paradigm. By training a single model jointly on multiple diverse datasets, we not only simplified the deployment pipeline but also demonstrably improved the generalization capabilities of various models on full-resolution data. This stands in contrast to previous generalization efforts, which often struggle with domain gaps. Finally, we presented a universally effective composite loss function that provides a significant performance uplift across all tested architectures, pushing the benchmarks for the field into a new era. We believe that our combined contributions—the PanTiny model, the all-in-one paradigm, and the powerful loss function—offer a more sustainable and practical path forward for future pan-sharpening research.
@@ A.2 The “One-Epoch Generalization” Illusion
To test the hypothesis of overfitting in conventional generalization studies, we conducted a surprising experiment: we trained several of our intermediate models for only one epoch on the WV2 dataset and then evaluated their performance on the unseen WV3 and GF2 datasets. The results, shown in Table 10, are striking. When tested on GF2, our one-epoch trained ‘M4 (Channel Attn)‘ model achieves a PSNR of 39.13. This result is comparable to or even surpasses the performance of fully-trained models from dedicated generalization papers, such as DDIF (Chen et al. 2022), which reports a PSNR of 37.77 on GF2 after being fully trained on WV2 (see Table 9). This suggests that the hundreds of additional training epochs in those works contribute little to true generalization, instead primarily reinforcing the model’s bias towards the source dataset. This finding strongly motivates a shift away from the “train-on-one, test-on-many” methodology. Any reviewer can easily verify this conclusion with a personal computer in under 5 minutes using our provided codebase, if they already have the datasets.
@@ A.3 The Overfitting Trap of Separate Training
Further evidence against the separate training paradigm comes from analyzing the cross-domain performance of our own model when fully trained on a single dataset. Table 11 shows the results of training ‘PanTiny (Big)’ to convergence on one source dataset and testing on all three. For instance, the model trained on WV2 achieves an excellent 42.16 PSNR on its own test set, but its performance plummets to 21.76 on WV3 and 33.92 on GF2. This performance is substantially worse than the one-epoch results, proving that prolonged training on a single dataset actively harms its ability to generalize by causing it to overfit to the source domain’s specific characteristics.
Table 11: Cross-domain performance of ‘PanTiny (Big)’ when trained separately on a single source dataset. The drastic performance drop on target datasets highlights the overfitting issue inherent in this paradigm.
@@ A.4 Failure Case: Generalization to Jilin-1 Dataset
To push the boundaries of generalization, we tested our all-in-one trained models on the Jilin-1 dataset, which was completely unseen during training. As shown in Table 13, the performance of all models is poor, indicating that even our robust ‘all-in-one’ paradigm has its limits when faced with a significant domain shift. Interestingly, PSCINN, which performed poorly on the training datasets, shows relatively better (though still low) performance here, possibly due to its different architectural inductive biases. This experiment reinforces our central thesis: true generalization in pan-sharpening is a data problem, and robust performance requires training on diverse, representative datasets.
Table 12: Full ablation study on loss function components and weights. The top part shows a broad search on our ‘PanTiny (Big)’ model, while the bottom part shows a fine-grained search on the ‘PanTiny (Small)’ model to accelerate experiments. Our proposed combination (1.5, 4.0, 1.5) provides the best overall performance. Best results are in bold, second-best are underlined.
@@
Table 13: Zero-shot generalization performance on the unseen Jilin-1 dataset. All models were trained under the ‘all-in-one’ paradigm. The best results are in bold and the second results are marked with underline.
Model Jilin-1
[…]
PNN 22.16 0.6000 0.1286
PanNet 22.82 0.6255 0.0911
PanFlow 22.14 0.5641 0.0861
MSDCNN 21.73 0.5988 0.1321
PSCINN 27.90 0.8319 0.0812
Ours (PanTiny Big) 23.10 0.5694 0.0884
Appendix B Detailed Ablation on the Composite Loss
B.1 The Overlooked Potential of Loss Functions
B.2 Our Systematic Two-Stage Search for the Optimal Loss
In the first stage, we conducted a broad search using our ‘PanTiny (Big)’ model to understand the general behavior of different loss component weightings. We tested balanced configurations like (1,1,1) as well as configurations focusing on each individual component. This initial exploration yielded a crucial insight: combinations with a high weight on the SSIM component, such as (1,3,1), consistently outperformed others.
Guided by this finding, we initiated a second, more fine-grained search stage. To accelerate experimentation, we used our lighter ‘PanTiny (Small)’ model and focused exclusively on high-SSIM weight combinations. This meticulous process allowed us to identify the ‘(1.5, 4.0, 1.5)’ configuration as the most robust and highest-performing combination. This discovery is not just a set of tuned hyperparameters; it represents a universally applicable principle that can elevate the entire field. By applying this composite loss, we have unlocked a new tier of performance, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era for a wide range of models.
@@ C.1 Downsampling Strategy
A common strategy in image restoration is to use a U-Net-like architecture with multiple downsampling stages to capture multi-scale features. We investigated this by creating variants of our base model (‘PanTiny(Small)’) with 0, 2, and 4 downsampling levels, using a basic L1 loss for a fair architectural comparison. As shown in Table 14, we found that increasing the downsampling levels led to a significant increase in parameters and a decrease in overall performance. The 0-level model (no downsampling) performed the best, indicating that for pan-sharpening, maintaining the full feature resolution is more effective. This led us to adopt a flat, single-scale architecture for PanTiny.
@@
C.2 Investigating the ‘DeepFusion’ Module
In our main paper, we noted that Pan-Mamba’s performance degrades significantly in the ‘all-in-one’ setting. We hypothesized this was due to its complex ‘DeepFusion’ module overfitting to single-dataset characteristics. To verify this, we integrated the ‘DeepFusion’ block into our ‘m6’ experimental model. As shown in Table 15, not only does the ‘DeepFusion’ block increase parameter count, but it also consistently underperforms compared to simpler fusion mechanisms like our ‘Enhanced Conv’ (from the main paper’s ablation) or even basic ‘Gated Conv’ and ‘Channel Attention’. Furthermore, increasing the depth of the ‘DeepFusion’ block from 2 to 5 layers leads to a further drop in performance. This provides strong evidence that such complex fusion modules, while effective for a single dataset, are detrimental to generalization in the ‘all-in-one‘ paradigm.
Table 15: Ablation on the ‘DeepFusion‘ module using our ‘m6‘ variant. Complex, deep fusion strategies underperform simpler ones in the multi-dataset setting.
@@ C.2 Investigating the ‘DeepFusion’ Module
In our architectural exploration, we also compared single-encoder and dual-encoder designs. Our ‘m5’ model variant features a dual-encoder architecture, while ‘m6’ uses a single encoder. Table 16 presents a controlled comparison where both models use a channel attention fusion mechanism. The ‘m6’ model, despite having significantly fewer parameters (64.3K vs. 118.5K), consistently outperforms the larger dual-encoder ‘m5’ model. This result was pivotal, leading us to abandon the more complex dual-encoder structure. We concluded that allocating parameters towards a more effective fusion and body in a single-encoder framework provides a better performance-efficiency trade-off, which became a core principle in designing the final ‘PanTiny’ model.
Table 16: Comparison between our single-encoder (‘m6’) and dual-encoder (‘m5’) experimental models. The single-encoder design achieves superior performance with fewer parameters.
@@ C.4 Full Ablation Results for Final Model Components
The main paper presented condensed versions of our final fusion and refinement ablations for brevity. Here, we provide the complete tables with all metrics (Table 17 and Table 18). These results reinforce our conclusion that for ‘PanTiny’, simple and well-chosen convolutional blocks outperform more complex alternatives in the multi-dataset setting, providing the best balance of parameter efficiency and performance.
Appendix D Additional Visual Results
Figure 6: Visual comparison on the WorldView-2 (WV2) dataset.
Figure 7: Visual comparison on the GaoFen-2 (GF2) dataset.
Table 17: Full ablation results for the fusion module in the final ‘PanTiny’ architecture. Our “Enhanced Conv” provides the best overall trade-off.
@@ Appendix D Additional Visual Results
Table 18: Full ablation results for the refinement module in the final ‘PanTiny’ architecture. A simple convolution is most effective.
@@ Appendix E Codebase and Reproducibility
To ensure full reproducibility and facilitate future research, we provide a comprehensive and easy-to-use codebase. Our framework is built around a unified experiment runner that leverages a hierarchical YAML configuration system. This allows researchers to define a base configuration and then specify a series of experiments that inherit and override these settings, enabling efficient and organized ablation studies. For more details, please refer to the ‘README.md’ on github.

### v2

@@ 5 Conclusion
In this paper, we challenged the prevailing “bigger is better” paradigm in pan-sharpening by focusing on efficiency, generalization, and principled engineering. We introduced PanTiny, a lightweight and highly efficient model, and demonstrated that its carefully considered architecture achieves a superior balance of performance and computational cost compared to larger, more complex SOTA models. Our most significant contribution is the novel “all-in-one” training paradigm. By training a single model jointly on multiple diverse datasets, we not only simplified the deployment pipeline but also demonstrably improved the generalization capabilities of various models on real-world, full-resolution data. This stands in contrast to previous generalization efforts, which often struggle with domain gaps.Finally, we presented a universally effective composite loss function that provides a significant performance uplift across all tested architectures, pushing the benchmarks for the field into a new era. We believe that our combined contributions—the PanTiny model, the all-in-one paradigm, and the powerful loss function—offer a more sustainable and practical path forward for future pan-sharpening research. Our code and pre-trained models will be made publicly available in the appendixs.
@@ A.2 The “One-Epoch Generalization” Illusion
To test the hypothesis of overfitting in conventional generalization studies, we conducted a surprising experiment: we trained several of our intermediate models for only one epoch on the WV2 dataset and then evaluated their performance on the unseen WV3 and GF2 datasets. The results, shown in Table 10, are striking. When tested on GF2, our one-epoch trained “M4 (Channel Attn)” model achieves a PSNR of 39.13. This result is comparable to or even surpasses the performance of fully-trained models from dedicated generalization papers, such as DDIF (Chen et al. 2022), which reports a PSNR of 37.77 on GF2 after being fully trained on WV2 (see Table 9). This suggests that the hundreds of additional training epochs in those works contribute little to true generalization, instead primarily reinforcing the model’s bias towards the source dataset. This finding strongly motivates a shift away from the “train-on-one, test-on-many” methodology. Any reviewer can easily verify this conclusion with a personal computer in under 5 minutes using our provided codebase, if they already have the datasets.
@@ A.3 The Overfitting Trap of Separate Training
Further evidence against the separate training paradigm comes from analyzing the cross-domain performance of our own model when fully trained on a single dataset. Table 11 shows the results of training “PanTiny (Big)” to convergence on one source dataset and testing on all three. For instance, the model trained on WV2 achieves an excellent 42.16 PSNR on its own test set, but its performance plummets to 21.76 on WV3 and 33.92 on GF2. This performance is substantially worse than the one-epoch results, proving that prolonged training on a single dataset actively harms its ability to generalize by causing it to overfit to the source domain’s specific characteristics.
@@
Table 11: Cross-domain performance of “PanTiny (Big)” when trained separately on a single source dataset. The drastic performance drop on target datasets highlights the overfitting issue inherent in this paradigm.
@@ A.4 Failure Case: Generalization to Jilin-1 Dataset
To push the boundaries of generalization, we tested our all-in-one trained models on the Jilin-1 dataset, which was completely unseen during training. As shown in Table 12, the performance of all models is poor, indicating that even our robust “all-in-one” paradigm has its limits when faced with a significant domain shift. Interestingly, PSCINN, which performed poorly on the training datasets, shows relatively better (though still low) performance here, possibly due to its different architectural inductive biases. This experiment reinforces our central thesis: true generalization in pan-sharpening is a data problem, and robust performance requires training on diverse, representative datasets.
Model Jilin-1
[…]
PNN 22.16 0.6000 0.1286
PanNet 22.82 0.6255 0.0911
PanFlow 22.14 0.5641 0.0861
MSDCNN 21.73 0.5988 0.1321
PSCINN 27.90 0.8319 0.0812
Ours (PanTiny Big) 23.10 0.5694 0.0884
Table 12: Zero-shot generalization performance on the unseen Jilin-1 dataset. All models were trained under the “all-in-one” paradigm. The best results are in bold and the second results are marked with underline.
Appendix B Detailed Ablation on the Composite Loss
B.1 The Overlooked Potential of Loss Functions
B.2 Our Systematic Two-Stage Search for the Optimal Loss
In the first stage, we conducted a broad search using our “PanTiny (Big)” model to understand the general behavior of different loss component weightings. We tested balanced configurations like (1,1,1) as well as configurations focusing on each individual component. This initial exploration yielded a crucial insight: combinations with a high weight on the SSIM component, such as (1,3,1), consistently outperformed others.
Guided by this finding, we initiated a second, more fine-grained search stage. To accelerate experimentation, we used our lighter “PanTiny (Small)” model and focused exclusively on high-SSIM weight combinations. This meticulous process allowed us to identify the “(1.5, 4.0, 1.5)” configuration as the most robust and highest-performing combination. This discovery is not just a set of tuned hyperparameters; it represents a universally applicable principle that can elevate the entire field. By applying this composite loss, we have unlocked a new tier of performance, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era for a wide range of models.
@@ B.2 Our Systematic Two-Stage Search for the Optimal Loss
Table 13: Full ablation study on loss function components and weights. The top part shows a broad search on our “PanTiny (Big)” model, while the bottom part shows a fine-grained search on the “PanTiny (Small)” model to accelerate experiments. Our proposed combination (1.5, 4.0, 1.5) provides the best overall performance. Best results are in bold, second-best are underlined.
@@ C.1 Downsampling Strategy
A common strategy in image restoration is to use a U-Net-like architecture with multiple downsampling stages to capture multi-scale features. We investigated this by creating variants of our base model (“PanTiny(Small)”) with 0, 2, and 4 downsampling levels, using a basic L1 loss for a fair architectural comparison. As shown in Table 14, we found that increasing the downsampling levels led to a significant increase in parameters and a decrease in overall performance. The 0-level model (no downsampling) performed the best, indicating that for pan-sharpening, maintaining the full feature resolution is more effective. This led us to adopt a flat, single-scale architecture for PanTiny.
@@
C.2 Investigating the “DeepFusion” Module
In our main paper, we noted that Pan-Mamba’s performance degrades significantly in the “all-in-one” setting. We hypothesized this was due to its complex “DeepFusion” module overfitting to single-dataset characteristics. To verify this, we integrated the “DeepFusion” block into our “m6” experimental model. As shown in Table 15, not only does the “DeepFusion” block increase parameter count, but it also consistently underperforms compared to simpler fusion mechanisms like our “Enhanced Conv” (from the main paper’s ablation) or even basic “Gated Conv” and “Channel Attention”. Furthermore, increasing the depth of the “DeepFusion” block from 2 to 5 layers leads to a further drop in performance. This provides strong evidence that such complex fusion modules, while effective for a single dataset, are detrimental to generalization in the “all-in-one” paradigm.
@@ C.2 Investigating the “DeepFusion” Module
Table 15: Ablation on the “DeepFusion” module using our “m6” variant. Complex, deep fusion strategies underperform simpler ones in the multi-dataset setting.
@@
In our architectural exploration, we also compared single-encoder and dual-encoder designs. Our “m5” model variant features a dual-encoder architecture, while “m6” uses a single encoder. Table 16 presents a controlled comparison where both models use a channel attention fusion mechanism. The “m6” model, despite having significantly fewer parameters (64.3K vs. 118.5K), consistently outperforms the larger dual-encoder “m5” model. This result was pivotal, leading us to abandon the more complex dual-encoder structure. We concluded that allocating parameters towards a more effective fusion and body in a single-encoder framework provides a better performance-efficiency trade-off, which became a core principle in designing the final “PanTiny” model.
@@
Table 16: Comparison between our single-encoder (“m6”) and dual-encoder (“m5”) experimental models. The single-encoder design achieves superior performance with fewer parameters.
@@ C.4 Full Ablation Results for Final Model Components
The main paper presented condensed versions of our final fusion and refinement ablations for brevity. Here, we provide the complete tables with all metrics (Table 17 and Table 18). These results reinforce our conclusion that for “PanTiny”, simple and well-chosen convolutional blocks outperform more complex alternatives in the multi-dataset setting, providing the best balance of parameter efficiency and performance.
@@
Table 17: Full ablation results for the fusion module in the final “PanTiny” architecture. Our “Enhanced Conv” provides the best overall trade-off.
@@
Table 18: Full ablation results for the refinement module in the final “PanTiny” architecture. A simple convolution is most effective.
Appendix D Additional Visual Results
Figure 5: Visual comparison on the WorldView-2 (WV2) dataset.
Figure 6: Visual comparison on the GaoFen-2 (GF2) dataset.
@@ Appendix E Codebase and Reproducibility
To ensure full reproducibility and facilitate future research, we provide a comprehensive and easy-to-use codebase as a “code.tar.gz” archive in the supplementary materials. Our framework is built around a unified experiment runner that leverages a hierarchical YAML configuration system. This allows researchers to define a base configuration and then specify a series of experiments that inherit and override these settings, enabling efficient and organized ablation studies. For more details, please refer to the “README.md” file included in our supplementary materials.